Magnetism: Find the missing using the 1st right hand rule
e) 8osT
Answer #1

Ans:-

Given data

B= 0.5T, V= 2*10^5m/s, q= 1.6*10^-19C, F= ? θ= 30deg

FB = qVxB= qvBsinθ

FB = 1.6*10-19*2*10^5*0.5*sin30

FB = 8*10^-15N